OpenStudy (anonymous):

Suppose you want to build a deck around your pool with a uniform width on all sides. The pool measures 15 yards by 20 yards. You want to increase the area of the pool to 600 yd2. What is the width of the deck? I'm sure it's simple I just don't get what I'm supposed to do.

OpenStudy (anonymous):

(20 + 2X) * (15 + 2X) = 600 300 + 40X + 30X + 4X^2 = 600 300 + 70X + 4X^2 = 600 4X^2 + 70X - 300 = 0 The real root of this equation is X = 3.561 yards This is the width of the deck.

OpenStudy (phi):

it is not *that* simple. First step is draw a picture

OpenStudy (anonymous):

would've taken me ages to figure out lol thank you! :)

OpenStudy (anonymous):

no problem:)